# RocqStar: Leveraging Similarity-driven Retrieval and Agentic Systems for Rocq generation

### v1

@@ 2.2 Modeling
In our work, we formulate the problem as a self-supervised contrastive representation learning problem and train a self-attentive embedder model [18]. Given the dataset of pairs $\left[\operatorname{statement}_{i},\operatorname{proof}_{i}\right]$, and a similarity function, $\operatorname{f}(\operatorname{proof}_{i},\operatorname{proof}_{j})$, defined between two proofs, we try to learn such function $r$ (ranker), that takes corresponding Rocq statements as inputs, but behaves as close as possible to $f$. Given two statements, we learn to predict how similar their proofs shall be. In § 4, we evaluate the performance of the proposed model on the following task. Given a statement $s_{*}$ and a set of proven theorems, we want to choose $k$ premises and use them as context for generating a proof for $S$.
$\mathcal{T}=\left\{(p_{i},s_{i})\right\},\;\mathcal{S}$ =$=\left\{s_{i}\right\},\;r:\mathcal{T}\times\mathcal{S}\to\mathbb{R}$ $\operatorname{Top}_{k}(r,s_{*})$ $=\underset{(p_{i},s_{i})\in\mathcal{T}}{\arg\operatorname{top}_{k}}\,g(s_{i},s_{*})$
$\operatorname{Solve}(r,s_{*})$ $=\operatorname{Solve}\bigl{(}\operatorname{Top}_{k}(r,s_{*}),s_{*}\bigr{)}\in\{0,1\}$ $\mathcal{Q}(r)$ $=\mathbb{E}_{s_{*}\sim\mathcal{D}}\bigl{[}\operatorname{Solve}(r,s_{*})\bigr{]}$
Assume, without loss of generality, that by basic statement similarity we mean $BM25$-based similarity. As we have already shown in § 2, text similarity is a bad choice of $r$, which shows low correlation with the target function. However, it sets a strong baseline for our model. In practical applications, similar theorems occasionally have similar proofs. Accordingly, we have decided to fine-tune Microsoft’s 108-million-parameter encoder [7], originally pretrained on a combined corpus of programming and natural language texts. On a tiny extra test dataset, consisting of 50 theorems with corresponding hand-picked premises, raw CodeBert achieved an accuracy of 48%. This corresponds to roughly the same accuracy for ranking performed using the Jaccard-similarity metric on statements.
@@
We train the model using InfoNCE [20] loss. In particular, given the statement $s$, on the dataset post-processing stage, we compute distances to other samples. We then mark a pair as positive if the distance between two proofs is less than a threshold $\tau_{pos}$, and we mark it as negative if the distance is greater than $\tau_{neg}$. Given the hyperparameter $k_{\operatorname{neg}}$ and sets of positive and negative pairs $P^{+}_{s}$ and $P^{-}_{s}$ we compute a per-statement loss term $\mathcal{L}_{s}$ as follows:
$\mathcal{L}_{s}=-\log\frac{\exp\bigl{(}\varphi(z_{s},z_{p})/T\bigr{)}}{\exp\bigl{(}\varphi(z_{s},z_{p})/T\bigr{)}+\displaystyle\sum_{j=1}^{k_{\operatorname{neg}}}\exp\bigl{(}\varphi(z_{s},z_{n_{j}})/T\bigr{)}}\qquad\bigl{(}p\in P^{+}_{s},\;n_{j}\in P^{-}_{s}\bigr{)}$
where $\varphi$ is a cosine similarity between $\ell_{2}$-normalized embeddings of statements. Experiments on the $k_{\operatorname{neg}}$ hyperparameter in our case showed little fluctuation in the results; however, $k_{\operatorname{neg}}=4$ procured the smoothest convergence, which aligns well with research by Wu et al. [29].
@@
Given the particular shape of the sample distance distribution, during training we experienced the problem of the model converging too quickly on “easy” negatives — pairs, whose proofs (and typically their statements) are already far apart in the raw distance space. To keep informative gradients flowing, we add hard negative pairs; with some probability we treat a pair of statements as negative if $\tau_{\operatorname{hardneg}}\leqslant\operatorname{sim}(\text{proof}_{a},\text{proof}_{b})\leqslant\tau_{\operatorname{neg}}$. Introduction of negative samples helped to stabilize the training process; we have observed a less steep training curve and better generalization overall. Other training hyperparameters are listed in Appendix C.
@@ 3 Agentic System
To allow interaction between the agent and Rocq’s system, we develop a REST API server that provides a set of tools that are useful during the execution. We apply our domain knowledge and construct these tools to bring an agent-driven proving process as close as possible to a human-driven one. Example of allowed function calls include checking validity of proofs, retrieving the valid prefix of given proof, gathering additional information about available entities in the context, and interacting with the context via performing commands like Print ?a. to identify the type of an argument or Search ?exp. to search for defined terms by a pattern. Toolset is described in detail in Appendix D. Interaction with Rocq’s system is carried out through its language server, coq-lsp [6]. To conform with a commonly used Model Context Protocol (MCP) and allow seamless agent interaction with the environment through tools, we implement an MCP server that wraps the REST API server. In the provided tool set, the most important is a proof-checking tool. It not only returns the answer whether the proof in question is valid, but in case of an erroneous proof, returns the error itself, where in the proof it happened, and the valid prefix before the error along with the remaining goals after this prefix. Such functionality enables the agent to keep track of the current proof state and benefit from partial proof progress.
@@ 3.1 Agent Logic
The input to the agent is presented as a target theorem without a proof and a file where it was declared, see box 31 of Fig. 2. Agent’s pipeline is logically split into two main stages: planning and execution. In the planning phase, multiple language models rigorously work out the strategy for the further implementation. During execution agents follow the plan aiming to generate the correct proof.
Planning Stage  We use the idea of multi-agent debates to produce a plan of how the agent should prove the given theorem. Specifically, we make two LLMs argue with each other about the plan: one of the LLMs produces the initial plan and defends it (pro LLM), the other one makes arguments against this plan (con LLM), see box 32 of Fig. 2. After several rounds of debates, the whole message history is sent to the judge LLM, that decides who is the winner of the debate, and returns the final plan. With this procedure, we generate $k$ plans. We send them to the plan scoring LLM and prompt it to assign a numerical score to each plan (the higher the better). After that, we select $l$ plans with the highest score and send them to the Execution Stage, see box 33 of Fig. 2.
Execution Stage  For each of the selected plans, we run an executor agent that follows the strategy, iteratively invoking the tools from the provided tool set — proof checker, context–inspection queries, search commands, and so on, as atomic actions. By calling these tools, it interacts with the environment via the MCP server. We track how many erroneous proofs were checked in a row, and if this number is higher than a fixed threshold (we set the threshold to five during evaluation), we call a critic model to evaluate the current progress of the proof and find the deviations from the selected plan. After that, we retrieve theorems along with their proofs, whose top-level goals are similar to the currently remaining goal, according to the cosine similarity between their RocqStar-ranker embeddings. We prompt the LLM to explain which tactic sequences could be helpful to finish our proof. We gather the generated criticism and send it to the replanner LLM to refine the current plan along with similar proofs and their analysis. The replanner is a separate language model that revises the plan based on the critic’s feedback and the retrieved examples. The whole message history is sent back to the executor agent. During the execution of each plan, $n$ tool calls are allowed. If valid proof is not found after $n$ tool calls, we denote the plan as failed. In this case, we ask a plan failure summarizer LLM to generate a short explanation of why the strategy execution failed and what happened during it. Then this summarized explanation is sent to the new execution stage with the next selected plan. This procedure is repeated until the correct proof is found or there are no more strategies to execute.
@@ 4.1 Retrieval Mechanism
We integrate our retrieval mechanism as a ranker into CoqPilot and evaluate it on the IMM-300 dataset with different models under the hood. We compare the performance of our ranker with the baseline approach, which works in the following manner. Given a target theorem statement $s_{*}$ and a set of proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, it ranks theorems in a descending order of $J(s_{*},s_{i})$, where $J(s_{*},s_{i})$ is Jaccard-similarity index and $S_{s_{i}}$ is a set of tokens inside a statement. The statement is split into tokens by whitespaces, commas, etc. Jaccard-similarity index is semantically almost the same as the BM-25 metric and produces the same numerical results. For each theorem in the dataset, we take theorems within the same file, sort them using the ranker (Jaccard or RocqStar, respectively), take the $k$ most relevant ones ($k$ is equal to 7 in our experiments), and send a request to the model to generate the completion. Chosen theorems are being sent as a few-shot prompt. Generation for each theorem is requested 12 times. If the Rocq’s system accepts any of the proofs, the theorem is considered solved. The target metric in our evaluation is the ratio of solved theorems. The evaluation results are presented in Table 1.
[…]
Ranker Jaccard RocqStar Jaccard RocqStar Jaccard RocqStar
GPT-4o 48% $\pm$ 5% 51%$\pm$ 5% 18% $\pm$ 4% 25%$\pm$ 3% 11% $\pm$ 4% 11% $\pm$ 5%
Claude 3.5 58% $\pm$ 5% 61%$\pm$ 4% 28% $\pm$ 5% 36%$\pm$ 5% 16% $\pm$ 5% 21%$\pm$ 5%
@@
As can be seen from Table 1, our proposed RocqStar ranker outperforms the baseline Jaccard ranker on almost all experiments, showing reliable improvement. Most of the performance increase could be seen in the second group; we interpret these results as follows. For short theorems in the first group, the assumption that similar statements imply similar proofs often holds; therefore, both rankers perform close to one another. For complex theorems from the third group, it rarely happens that two theorems have significantly similar proofs, resulting in less advancement space for the model.
@@ 4.2 Agentic System
As shown in Table 2, our agentic system outperforms other benchmarked models inside the CoqPilot framework. The strongest model so far was Claude 3.5 Sonnet, which achieves 51% accuracy on the dataset, given 12 retries for each theorem. RocqStar agent achieves 60%, showing vigorous improvement. In terms of financial costs, we estimate a run of an agent on one theorem at 1.3 US dollars, compared to 0.25 US dollars for 12 requests to the pure Claude 3.5 Sonnet in CoqPilot.
@@ 4.3 Ablation study
Considering that software-verification tasks cannot be solved ad hoc, without explicit planning, we conduct an ablation study that measures how removing the Multi-Agent Debate (MAD) layer and reverting to single-pass planning affects the proportion of theorems successfully proved. In this experiment, we leave all other modules of the system unchanged, including the Plan Scoring LLM. We run two versions of an agent; the first generates plans via MAD, and the other generates plans by a single request to an LLM without further refinement. We evaluate both agents on the IMM-50 dataset and depict the results in Table 3. The same table depicts the performance of the Claude 3.5 Sonnet model on the IMM-50 dataset with 12 retries and RocqStar ranker; this result is provided for reference.
@@
Results in Table 3 confirm a consistent advantage for MAD across all three complexity groups, with the most significant improvement observed on harder theorems. This trend highlights the importance of MAD in solving composite multi-stage problems, such as complex proof construction. We attach an example of how MAD refines the execution plan and fixes it to manage to solve a previously unsolved theorem in Appendix E.

### v2

@@ 2.2 Modeling
In our work, we formulate the problem as a self-supervised contrastive representation learning task and train a self-attentive embedder model [19]. Given a dataset $\mathcal{T}=\{(s_{i},p_{i})\}$, where $s_{i}$ is a Rocq statement and $p_{i}$ is its corresponding proof, and a similarity function $\operatorname{f}(\operatorname{proof}_{i},\operatorname{proof}_{j})$, defined between two proofs, we aim to learn a function
$r:\mathcal{S}\times\mathcal{S}\to\mathbb{R},$
which takes two statements $s_{i},s_{j}\in\mathcal{S}$ as inputs and outputs a score approximating the similarity of their respective proofs $p_{i},p_{j}$. In other words, $r(s_{i},s_{j})\approx f(p_{i},p_{j})$. This formulation allows the ranker $r$ to assign scores to candidate statements relative to a target statement, thereby guiding retrieval towards those whose proofs are most likely to be useful.
In §˜4, we evaluate the performance of the proposed model in the following task. Given a target statement $s_{*}$ and a set of proven theorems $\mathcal{T}$, we aim to select $k$ premises from $\mathcal{T}$ to be used as context for generating a proof of $s_{*}$. We take $k$ most relevant theorems, according to the ranker $r$.
$\operatorname{Top}_{k}(r,s_{*})=\underset{(s_{i},p_{i})\in\mathcal{T}}{\arg\operatorname{top}_{k}}\,r(s_{i},s_{*})$
We say that statement $s_{*}$ is solved with the use of the ranker $r$ if the generator $g$ produces a valid proof for $s_{*}$ given the premises selected by $r$.
$\operatorname{Solve}(s_{*},r,g)=\begin{cases}1,&g\bigl(\operatorname{Top}_{k}(r,s_{*}),\,s_{*}\bigr)\;\text{is a valid proof},\\ 0,&\text{otherwise}.\end{cases}$
Finally, the quality of the ranker is estimated by the number of theorems in the evaluation set that can be solved using $r$ in combination with a given generator $g$.
One of the difficulties encountered during training is the U-shaped distribution of proof distances over random theorem pairs. Many short proofs were extremely similar to one another, while others were mostly too far apart, leaving a mid-range gap that hindered training. To mitigate this, we slightly modify the definition of $\mathrm{proof\_distance}(\cdot)$ introduced in §˜2 incorporating an additional similarity term and noise for robustness.
$\mathrm{proof\_distance}(p_{i},p_{j})$ $=\alpha\,D_{L}(p_{i},p_{j})+(1-\alpha)\,D_{J}(p_{i},p_{j})+\gamma$
$D_{J}(p_{i},p_{j})$ $=1-\frac{|p_{i}\cap p_{j}|}{|p_{i}\cup p_{j}|}$
The coefficient $\alpha=0.7$ was chosen heuristically based on the distribution plot and yielded the best performance in experiments. The noise $\gamma$ is taken from $\mathcal{U}(-1e-3,+1e-3)$.
As we have already shown in §˜2, statement similarity is a poor choice of $r$, as it shows a low correlation with the target function $\mathrm{proof\_distance}$. However, it still provides a strong baseline: in practice, similar theorems occasionally have similar proofs. Accordingly, our approach builds upon statement encoders, fine-tuning them to better align with the underlying proof structure. We fine-tune Microsoft’s 108-million-parameter encoder CodeBert [8], originally pretrained on a combined corpus of programming and natural language texts. We also experimented with gte-modernbert-base444gte-modernbert-base: https://huggingface.co/Alibaba-NLP/gte-modernbert-base as the base model, but it did not yield notable improvements. As shown in §˜4.1, the encoder without post-training performs on par with the Jaccard-similarity baseline, indicating that the unadapted model relies primarily on surface-level syntactic similarity rather than proof-related semantics. Our goal, therefore, is to adapt the encoder to capture this deeper semantic relation.
@@
To achieve this, we train the model using the InfoNCE [21] loss. In our setting, the distribution of proof distances is imbalanced even after normalization. InfoNCE naturally handles this case by contrasting a limited number of positives against a set of negatives, ensuring that informative gradients are maintained. In particular, given a statement $s$, during dataset preprocessing we compute distances to other samples. We then mark a pair as positive if the distance between their proofs is less than a threshold $\tau_{pos}$, and mark it as negative if it is greater than $\tau_{neg}$. Given the hyperparameter $k_{\operatorname{neg}}$ and sets of positive and negative pairs $P^{+}_{s}$ and $P^{-}_{s}$, we compute a per-statement loss term $\mathcal{L}_{s}$ as follows:
$\mathcal{L}_{s}=-\log\frac{\exp\bigl(\cos(z_{s},z_{p})/T\bigr)}{\exp\bigl(\cos(z_{s},z_{p})/T\bigr)+\displaystyle\sum_{j=1}^{k_{\operatorname{neg}}}\exp\bigl(\cos(z_{s},z_{n_{j}})/T\bigr)}$
where $\cos$ is a cosine similarity between $\ell_{2}$-normalized embeddings of statements, and $p\in P^{+}_{s}$, $n_{j}\in P^{-}_{s}$. On average, we observed smoother convergence for higher values of $k_{\operatorname{neg}}$, which is consistent with findings by Chen et al. [5]. However, due to hardware limitations, we selected $k_{\operatorname{neg}}=100$ as a practical trade-off between convergence stability and computational cost.
@@
Despite the adjustment of $\mathrm{proof\_distance}(\cdot)$, during training we experienced the problem of the model converging too quickly on “easy” negatives — pairs, whose proofs (and typically their statements) are already far apart in the raw distance space. To keep informative gradients flowing, we add hard negative pairs; with some probability we treat a pair of statements as negative if $\tau_{\operatorname{hardneg}}\leqslant\operatorname{sim}(\text{proof}_{a},\text{proof}_{b})\leqslant\tau_{\operatorname{neg}}$. Introduction of negative samples helped to stabilize the training process; we have observed a less steep training curve and better generalization overall. Other training hyperparameters are listed in Appendix C.
@@ 3 Agentic System
To allow interaction between the agent and Rocq’s system, we develop a REST API server that provides a set of tools that are useful during the execution. We apply our domain knowledge and construct these tools to bring an agent-driven proving process as close as possible to a human-driven one. Examples of allowed function calls include checking validity of proofs, retrieving the valid prefix of given proof, gathering additional information about available entities in the context, and interacting with the context via performing commands like Print ?a to identify the type of an argument or Search ?exp to search for defined terms by a pattern. Toolset is described in detail in Appendix D. Interaction with Rocq’s system is carried out through its language server, coq-lsp [7]. To conform with a commonly used Model Context Protocol (MCP) and allow seamless agent interaction with the environment through tools, we implement an MCP server that wraps the REST API server. Among the provided tools, the most important is the proof-checking tool. It not only verifies whether a proof is valid but, in case of an error, returns detailed diagnostic information: the error message, its exact location, the valid prefix preceding the error, and the remaining goals after that prefix. This functionality allows the agent to maintain awareness of the current proof state and leverage partial proof progress.
@@ 3.1 Agent Logic
The input to the agent is presented as a target theorem without a proof and a file where it was declared, see box 1 of Fig.˜2. Agent’s pipeline is logically split into two main stages: planning and execution. In the planning phase, multiple language models rigorously work out the strategy for the further implementation. During execution agents follow the plan aiming to generate the correct proof.
Planning Stage  We employ the idea of multi-agent debates to generate a strategy for proving the given theorem. Specifically, two LLMs engage in a discussion: one proposes and defends an initial plan (pro LLM), while the other critiques it (con LLM); see box 2 in Fig.˜2. After several debate rounds, the entire message history is passed to a judge LLM, which determines the winner and produces the final plan. Using this procedure, we generate $k$ candidate strategies. These are then evaluated by a plan scoring LLM that assigns each a numerical score (the higher, the better). Finally, the top-$l$ plans are selected and forwarded to the Execution Stage; see box 3 of Fig.˜2.
Execution Stage  For each of the selected plans, we run an executor agent that follows it step by step, invoking tools from the provided tool set — proof checker, context–inspection queries, search commands, and others, as atomic actions. Through these tool calls, the agent interacts with the environment via the MCP server. In addition to this iterative execution, we employ a reflection mechanism that monitors the progress of the proof and adjusts the strategy when necessary. We track how many consecutive erroneous proof attempts occur, and once this number exceeds a predefined threshold (set to five during evaluation), a critic model is called to assess the current proof state and identify deviations from the intended strategy. After that, we retrieve theorems along with their proofs, whose top-level goals are similar to the currently remaining goal, according to the cosine similarity between their RocqStar-ranker embeddings. We prompt the LLM to explain which tactic sequences could be helpful to finish our proof. We gather the generated criticism and send it to the replanner LLM to refine the current plan along with similar proofs and their analysis. The replanner is a separate language model that revises the plan based on the critic’s feedback and the retrieved examples. The whole message history is sent back to the executor agent. During the execution of each plan, $n$ tool calls are allowed. If valid proof is not found after $n$ tool calls, we denote the plan as failed. In this case, we ask a plan failure summarizer LLM to generate a short explanation of why the strategy execution failed and what happened during it. Then this summarized explanation is sent to the new execution stage with the next selected plan. This procedure is repeated until the correct proof is found or there are no more strategies to execute.
@@ 4.1 Retrieval Mechanism
We integrate our retrieval mechanism as a ranker into CoqPilot and evaluate it on the IMM-300 dataset with different models under the hood. To assess its performance, we compare our approach against two baselines: (i) an untrained embedder model (we use gte-modernbert-base, with codebert-base yielding comparable results), and (ii) a lexical similarity baseline based on the Jaccard index. In the latter, given a target theorem statement $s_{*}$ and a set of proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, it ranks the theorems in descending order of $J(s_{*},s_{i})$, where $J(s_{*},s_{i})$ is the Jaccard-similarity index and $S_{s_{i}}$ is a set of tokens inside a statement. The statement is split into tokens by whitespaces, commas, etc. Jaccard-similarity index is semantically almost the same as the BM-25 metric and produces the same numerical results. For each theorem in the dataset, we take theorems within the same file, sort them using the ranker (Jaccard, ModernBert, or RocqStar, respectively), take the $k$ most relevant ones ($k$ is equal to 7 in our experiments) and send a request to the model to generate the completion. The chosen theorems are being sent as a few-shot prompt. Generation for each theorem is requested 12 times. If the Rocq’s system accepts any of the proofs, the theorem is considered solved. The target metric in our evaluation is the ratio of solved theorems. The evaluation results are presented in Table 1. The reported values denote mean success rates, and the $\pm$ intervals correspond to the standard deviation across three independent runs.
Group $\mathbf{\leqslant 4}$ $\mathbf{5\,-\,8}$ $\mathbf{9\,-\,20}$
Ranker Jaccard ModernBert RocqStar Jaccard ModernBert RocqStar Jaccard ModernBert RocqStar
GPT-4o 48% $\pm$ 5% 44% $\pm$ 4% 51% $\pm$ 5% 18% $\pm$ 4% 21% $\pm$ 5% 25% $\pm$ 3% 11% $\pm$ 4% 8% $\pm$ 4% 14% $\pm$ 5%
Claude 3.5 58% $\pm$ 5% 57% $\pm$ 3% 61% $\pm$ 4% 28% $\pm$ 5% 30% $\pm$ 3% 36% $\pm$ 5% 16% $\pm$ 5% 16% $\pm$ 5% 21% $\pm$ 5%
@@
As can be seen from Table 1, our RocqStar ranker consistently outperforms both the Jaccard baseline and the untrained ModernBert encoder, demonstrating reliable gains across all evaluation groups. Most of the performance increase could be seen in the second group; we interpret these results as follows. For short theorems in the first group, the assumption that similar statements imply similar proofs often holds; therefore, all rankers perform comparably. For complex theorems from the third group, it rarely happens that two theorems have significantly similar proofs, resulting in less advancement space for the model.
@@ 4.2 Agentic System
As shown in Table 2, our agentic system outperforms other benchmarked models inside the CoqPilot framework. The strongest model so far was Claude 3.5 Sonnet, which achieves 51% accuracy on the dataset, given 12 retries for each theorem. RocqStar agent achieves 60%, showing vigorous improvement. In terms of financial costs, we estimate a run of an agent on one theorem at 1.3 US dollars, compared to 0.25 US dollars for 12 requests to the pure Claude 3.5 Sonnet in CoqPilot. Along with five language models invocated through the CoqPilot framework, we have compared our solution to other Rocq generation approaches, such as Tactician and Rango. On our IMM-300 dataset both solutions showed a result comparable to CoqPilot with OpenAI GPT-4o as the generator model.
@@ 4.3 Ablation study
We conduct an ablation study to analyze the contribution of individual components of the agentic pipeline to the overall success rate. In particular, we investigate the effects of removing (1) the Multi-Agent Debate (MAD) layer responsible for iterative plan refinement, (2) the Planning stage entirely, (3) the RocqStar retrieval module, and (4) the Reflection mechanism responsible for forced retrieval, criticism, and replanning. All experiments are performed on the IMM-50 dataset, with all other system components kept unchanged. The results are summarized in Table 3.
@@
Agent w/o Planning 86% 50% 17% 58%
Agent w/o RocqStar retrieval 86% 50% 33% 62%
Agent w/o Reflection 73% 44% 8% 48%
@@
Planning  Considering that software-verification tasks cannot be solved ad hoc, without explicit planning, we measure how removing the MAD layer and reverting to single-pass planning affects the proportion of successfully proved theorems. We run two versions of the agent: one generates plans via MAD, and the other produces a single plan in one LLM call without further refinement. Additionally, we include a configuration with the Planning stage entirely disabled, where the executor immediately attempts to construct a proof without any plan. This comparison shows that an agent without planning performs nearly identically, or slightly worse, than one guided by a poor single-pass plan, suggesting that a suboptimal plan does not provide advantage. In contrast, multi-step planning via MAD yields a consistent improvement across all groups, confirming the importance of structured plan refinement. An example of how MAD repairs a previously unsuccessful plan is presented in Appendix E.
RocqStar retrieval  To further justify the improvements of the agentic pipeline, we evaluate the agent with its retrieval component replaced by the baseline Jaccard index. The observed drop in performance compared to the full pipeline highlights the importance of the RocqStar retrieval approach, while the gain over the Claude 3.5 Sonnet model demonstrates the effectiveness of the proposed agentic design.
Reflection  Finally, we disable the reflection mechanism that triggers forced retrieval and replanning after failed attempts. Without reflection, the agent loses its ability to recover from early mistakes, leading to a noticeable degradation of performance, especially on longer and more complex proofs. The result shows that reflection complements planning by enabling recovery from failed reasoning trajectories.